OpenStudy (anonymous):

12) Fill in the blanks for the angle measures associated with a full circle: _______ radians = ______°

jimthompson5910 (jim_thompson5910):

Think of a circle with radius 1 This circle has a circumference of C = 2*pi*r C = 2*pi*1 C = 2pi So 360 degrees, aka a full revolution of a complete circle, is equal to 2pi radians. You can cut each part in half to get 2pi radians = 360 degrees pi radians = 180 degrees